E6-18 Analyzing Multiproduct CVP [LO 6-6]

Tiago makes three models of camera lens. Its product mix and contribution margin per unit follow:

Percentage of Unit sales Contribution Margin per unit
Lens A 28 % $ 35
Lens B 41 27
Lens C 31 40


Required:
1.
Determine the weighted-average contribution margin per unit. (Round your intermediate calculations and final answer to 2 decimal places.)

Weighted Average CM per unit




2. Determine the number of units of each product that Tiago must sell to break even if fixed costs are $192,000. (Round intermediate calculations and final answers to the nearest whole number.)

Units of A Units
Units of B Units
Units of C Units




3. Determine how many units of each product must be sold to generate a profit of $78,000. (Round intermediate calculations and final answers to the nearest whole number.)

Units of A Units
Units of B Units
Units of C Units
Answer #1

1) Weighted average contribution margin per unit = (35*28%+27*41%+40*31%) = 33.27

2) Break even point (unit) = 192000/33.27 = 5771 Units

Units of A (5771*28% 1616 Units
Units of B (5771*41% 2366 Units
Units of C (5771*31%) 1789 Units

3) Target unit (unit) = 192000+78000/33.27 = 8115 Units

Units of A (8115*28% 2272 Units
Units of B (8115*41% 3327 Units
Units of C (8115*31%) 2516 Units